OpenStudy (anonymous):

solve for 0 < or equal to x<2pie, give answers to the nearest hundredth of a radian if necessary. Sin x + Cos x = 0

OpenStudy (anonymous):

x= (4n-1) pi/4... Use sinx + cosx = \[\sqrt{2} \sin (\pi/4 + x) \] For the given interval of 'x' check which values of 'n' give you the desired solution set... Hope it makes sense :-)

OpenStudy (anonymous):

ok :) thankyou

OpenStudy (anonymous):

My pleasure :-)